### v1

# Dense 3D Displacement Estimation for Landslide Monitoring via Fusion of TLS Point Clouds and Embedded RGB Images

Abstract
Landslide monitoring is essential for understanding geohazards and mitigating associated risks. However, existing point cloud-based methods typically rely on either geometric or radiometric information and often yield sparse or non-3D displacement estimates. In this paper, we propose a hierarchical partition-based coarse-to-fine approach that fuses 3D point clouds and co-registered RGB images to estimate dense 3D displacement vector fields. We construct patch-level matches using both 3D geometry and 2D image features. These matches are refined via geometric consistency checks, followed by rigid transformation estimation per match. Experimental results on two real-world landslide datasets demonstrate that our method produces 3D displacement estimates with high spatial coverage (79% and 97%) and high accuracy. Deviations in displacement magnitude with respect to external measurements (total station or GNSS observations) are 0.15 m and 0.25 m on the two datasets, respectively, and only 0.07 m and 0.20 m compared to manually derived references. These values are below the average scan resolutions (0.08 m and 0.30 m). Our method outperforms the state-of-the-art method F2S3 in spatial coverage while maintaining comparable accuracy. Our approach offers a practical and adaptable solution for TLS-based landslide monitoring and is extensible to other types of point clouds and monitoring tasks. Our example data and source code are publicly available at github.com/zhaoyiww/fusion4landslide.
keywords:
Landslide monitoring , terrestrial laser scanning , RGB image , coarse-to-fine , displacement vector field
journal: Journal of  Templates
1 Introduction
Landslides occur worldwide at local and regional scales, often posing significant geological hazards, particularly when they happen near human infrastructure and activities. These hazards can be triggered by natural processes (e.g., earthquake shaking, erosion, and rainfall), human activities (e.g., deforestation, construction, and excavation), or a combination of these factors (Zhao and Lu, 2018; Casagli et al., 2023; Albanwan et al., 2024). Their consequences can be catastrophic; for example, 55,997 fatalities and widespread infrastructure damage were recorded in 4,862 landslide events worldwide between 2004 and 2016 (Froude and Petley, 2018). Given the severity of these impacts, monitoring landslide behavior and changes over time (i.e., landslide monitoring) is essential to risk mitigation, facilitating effective risk management, early warning, and prediction (Abellán et al., 2014; Casagli et al., 2017).
Landslide monitoring involves measuring and estimating displacements of landslides at specific points over time (i.e., across multiple epochs). Remote sensing techniques (RSTs) are indispensable for this task due to their non-contact nature, allowing the observation of geometric and/or radiometric changes without direct access to hazardous or inaccessible areas (Mantovani et al., 1996; Casagli et al., 2017). RSTs can be operated across multiple platforms, including spaceborne systems such as satellite imagery (Stumpf et al., 2014) and synthetic aperture radar (SAR) (Liu et al., 2021), airborne systems such as airborne LiDAR (Ventura et al., 2011) and unmanned aerial vehicle (UAV) photogrammetry (Lucieer et al., 2014), and ground-based systems such as ground-based SAR (Casagli et al., 2010) and terrestrial laser scanning (TLS) (Pesci et al., 2011). These tools enhance landslide monitoring by offering different capabilities in range, accuracy, and spatial and temporal resolution. Among them, TLS stands out for its capability to rapidly acquire dense 3D point clouds (e.g., thousands of points per second) with high spatial resolution and measurement accuracy (Jaboyedoff et al., 2012; Travelletti et al., 2014; Casagli et al., 2023).
To estimate displacements from multi-epoch point clouds (e.g., captured by TLS scanners), many approaches have been proposed and validated in different scenarios (Lindenbergh and Pietrzyk, 2015; Casagli et al., 2023). These methods can be broadly categorized based on the type of input information, and the nature of their outputs, namely displacement magnitudes along a predefined direction or full 3D displacement vectors. Traditional methods such as C2C (Girardeau-Montaut et al., 2005), C2M (Cignoni et al., 1998), and original M3C2 (Lague et al., 2013) rely on distances computed in Euclidean space to establish point-to-(point; approximated surface; averaged points) correspondences. While simple and widely used, these methods are sensitive to point density variation and surface roughness. More recently, deep learning-based approaches have emerged that establish point-wise correspondences in a high-dimensional feature space, with features extracted by pretrained neural networks. These methods exploit the strong generalizability of local patch-based networks, enabling deployment of models trained on annotated indoor datasets to unlabeled outdoor landslide scenes. While these approaches have shown promising results, only a few studies have explored this direction for point cloud-based landslide monitoring (Gojcic et al., 2020, 2021; Kenner et al., 2022). A more detailed description of methods of different categories is provided in Section 2.1.
On the one hand, most existing methods for estimating landslide displacements from 3D point clouds rely on a single type of information, either geometric or radiometric. While some studies (Jiang et al., 2022; Zhou et al., 2024) integrate TLS and UAV-based photogrammetry for monitoring, such fusion is primarily at the data level—aiming to generate denser point clouds—while further analysis still relies solely on a single modality, such as geometry. However, using one modality may fail to estimate accurate displacements in certain scenarios that could be recovered by the other. For instance, radiometry-based approaches have been shown to perform well in regions characterized by planar or repetitive geometric structures, where the presence of radiometric texture is evident (Wang et al., 2025b). Conversely, geometry-based methods have difficulty in achieving accurate estimation (Wang et al., 2025a).
On the other hand, many TLS scanners contain built-in, calibrated RGB cameras that produce high-resolution RGB images, along with manufacturer-provided parameters for co-registration with the point clouds. However, these images are primarily used for visualization (Balado et al., 2023) rather than for supporting displacement estimation. In our previous work (Wang et al., 2025a), we explored the feasibility of using radiometric information from these RGB images to establish correspondences for displacement estimation, and found that using radiometric data can complement geometry-based methods. Given these findings, we argue that leveraging both modalities is essential for estimating displacement vectors with high spatial coverage and accuracy in different unfavorable situations, e.g., geometric smoothness and high illumination changes.
To address the aforementioned issues, we propose a deep learning-based, cross-modal approach to estimating dense 3D displacements for TLS-based landslide monitoring. Our key innovation lies in first performing hierarchical partition-based coarse (patch) matching, followed by fine matching between each matched patch pair. This process considers both point clouds and RGB images. To mitigate unreliable correspondences computed purely in Euclidean space, we instead adopt feature-based correspondences that use deep features extracted by pre-trained deep learning models with strong generalization (Poiesi and Boscaini, 2021), inspired by the strategy in F2S3 (Gojcic et al., 2021). Furthermore, to avoid the limitations of uniform partitioning, which can disrupt object boundaries by ignoring underlying structures, and single-level partitioning, which may sacrifice spatial coverage to reduce the number of noisy displacement estimates, we adopt an advanced hierarchical partitioning method. This method partitions input point clouds into three levels of partitions, taking geometric and (optionally) radiometric features into account. By optimizing the matching pipeline and fusing 3D point clouds and 2D RGB information, our approach achieves displacement estimation accuracy comparable to the current state-of-the-art method (F2S3) while improving spatial coverage.
The remainder of the paper is organized as follows: Section 2 provides a literature review of point cloud-based displacement estimation, as well as 2D and 3D deep learning-based feature description methods. Section 3 elaborates the principle of the proposed method for estimating 3D displacement vectors for landslide monitoring. Section 4 illustrates study areas and datasets, data preprocessing, and evaluation. Section 5 presents the experimental results and discussion. Finally, Section 6 summarizes the paper and outlines future research directions.
2 Related work
2.1 Point cloud-based displacement estimation
Many methods have been proposed to estimate displacements from point clouds acquired at different epochs, To estimate displacements from point clouds, either from TLS scanners or from photogrammetric 3D reconstructions. Here we broadly classify these methods into five categories. Category I includes traditional methods such as C2C (Girardeau-Montaut et al., 2005), C2M (Cignoni et al., 1998), and original M3C2 (Lague et al., 2013), which estimate displacements by comparing spatial distances in Euclidean space, typically along surface normals. These methods operate via nearest neighbor (NN) searches or projection onto local reference surfaces, using strategies such as point-to-point, point-to-mesh, or averaged point-to-neighborhood comparisons. While effective in many cases, they are sensitive to variations in point density and surface roughness and often miss tangential displacements occurring along the surface. Several adaptations James et al. (2017); Zahs et al. (2022); Yang and Schwieger (2023) aim to improve correspondence quality or enhance the selection of informative directional estimates. Multi-directional M3C2 (Williams et al., 2021) improves flexibility by considering multiple measurement directions. However, without explicit point-wise correspondence tracking, these methods estimate displacement only along predefined direction(s) rather than recovering full 3D displacement vectors.
Category II entails “Piecewise ICP” and related alternatives (Teza et al., 2007; Friedli and Wieser, 2016; Wujanz et al., 2018), which uniformly partition point clouds into tiles and estimate per-tile displacements as the norms of translation vectors obtained through ICP algorithm (Besl and McKay, 1992; Chen and Medioni, 1992; Bergevin et al., 1996) or some of its variants. In this case, the per-tile correspondences are also established directly in the Euclidean space. Despite their effectiveness in certain scenarios, this class of methods produces yields biased estimates in instances where per-tile rigid body motion is violated or points within the tile are inadequate for accurate estimation.
Category III consists of methods that estimate 3D displacement vectors by establishing per-point correspondences using deep learning-based 3D point feature descriptors, such as F2S3 (Gojcic et al., 2020, 2021). Category IV relies on image-based representations of the acquired point clouds, typically through hillshade rendering, or directly using RGB images (Wang et al., 2025a) to establish correspondences in 2D image space. The estimations of these methods are typically based on image correlation, optical flow, or feature matching (Fey et al., 2015; Holst et al., 2021; Aati et al., 2022).
Category V encompasses methods developed for point clouds generated through photogrammetric 3D reconstruction rather than TLS. For instance, Wagner (2016) employs a total station equipped with a telescope camera to obtain 3D observations and RGB images simultaneously. Matching is then conducted on RGB images using cross-correlation methods. Lucks et al. (2024) applies SIFT (Lowe, 2004) to UAV images to identify 2D matches, which are then reconstructed into 3D displacements using photogrammetric techniques. Similarly, Ioli et al. (2023, 2024) perform dense 3D reconstruction from UAV data for glacier monitoring. They utilize deep-learning-based image matching algorithms for matching, leveraging deep-learning-based image matching to yield denser displacement vectors than earlier approaches. While these methods are relatively easy to implement, their accuracy heavily depends on reconstruction quality and they are often sensitive to large illumination variations and seasonal changes.
2.2 2D, 3D deep learning-based feature description
Detector-based feature description methods
Non-learning-based 2D pixel-wise matching algorithms, such as SIFT (Lowe, 2004), are constrained by their sensitivity to large illumination changes and viewpoint variations (Sun et al., 2021). To address these challenges, a number of studies have proposed the utilization of deep learning networks for pixel-wise matching (Dusmanu et al., 2019; Revaud et al., 2019; Luo et al., 2020). Inspired by these works, Bai et al. (2020) propose using joint feature detection and description for 3D point-wise matching. Huang et al. (2021) introduce a method that predicts keypoint saliency and point cloud overlap scores to enhance detection and description. This approach specifically addresses the challenge of aligning point clouds with low overlap. Although these detector-based methods have improved the performance of pixel-wise or point-wise matching, they often struggle when keypoint detection is ineffective. Consequently, inaccurate keypoint detection can also limit the performance of feature descriptors.
Detector-free feature description methods
Instead of explicitly detecting and describing keypoints and then matching them, detector-free methods integrate feature matching directly into the learning process. These methods typically employ a coarse-to-fine step to implement matching (Li et al., 2020; Zhou et al., 2021). More recently, LoFTR (Sun et al., 2021) and its variant Efficient LoFTR (Wang et al., 2024), introduce transformer-based architectures to improve the accuracy of 2D image matching. Inspired by the coarse-to-fine matching mechanism used for 2D image matching, CoFiNet (Yu et al., 2021) adopt a similar mechanism for the 3D point matching task. Building on this framework, GeoTransformer (Qin et al., 2022) incorporate transformer networks to further improve registration accuracy, particularly in challenging low-overlap scenarios.
Our proposed method is detector-free and introduces two key innovations over existing detector-free approaches: (i) Existing methods rely on fully convolutional networks that process entire point clouds, which limits their applicability to large-scale datasets. In contrast, our method leverages a local patch-based network (Poiesi and Boscaini, 2021) with strong generalizability. This design is especially suited for landslide monitoring, where point clouds may contain tens of millions of points and labeled training data are scarce. (ii) Existing methods often employ uniform subsampling to generate hierarchical keypoints or patches (Thomas et al., 2019). Instead, we adopt a hierarchical, feature-based partitioning strategy that aligns better with object boundaries, enabling more effective motion pattern estimation.
3 Methodology
Figure 1: Method overview. The proposed pipeline consists of five steps: (1) Point cloud tiles and the corresponding image tiles are generated. (2) Each paired source-target point cloud tile is partitioned into smaller patches. (3) Coarse matching establishes patch-wise correspondences / matches based on both 3D point cloud and 2D RGB information. (4) A refinement module evaluates match quality and discards low-quality patch matches based on predefined criteria. (5) For each pair of patch matches, a rigid transformation is estimated and applied to all points within the source patch. The transformed source points and associated displacement vectors are then aggregated to construct the final 3D DVF, optionally producing point-wise matches.
Given two TLS-acquired point clouds of a landslide at different epochs, denoted as the source ($\mathbf{P}$) and target ($\mathbf{Q}$) point clouds, along with their associated RGB images ($\mathbf{I}^{\mathrm{p}}$ and $\mathbf{I}^{\mathrm{q}}$) obtained from built-in calibrated cameras, our objective is to estimate a dense 3D displacement vector field (DVF) for landslide monitoring. Formally, by jointly leveraging both 3D geometry and 2D RGB information, we aim to compute, for each point $(x,y,z)$ in $\mathbf{P}$, a displacement vector $(\Delta X,\Delta Y,\Delta Z)$ that represents its estimated motion from $\mathbf{P}$ toward its most likely counterpart in $\mathbf{Q}$:
$\mathbf{V}=\mathcal{F}(\mathbf{P},\mathbf{Q},\mathbf{I}^{\mathrm{p}},\mathbf{I}^{\mathrm{q}}),$ (1)
where $\mathbf{V}:\mathbb{R}^{3}\ni(x,y,z)\mapsto(\Delta X,\Delta Y,\Delta Z)\in\mathbb{R}^{3}$ denotes the 3D DVF, and $\mathcal{F}$ is the function that estimates $\mathbf{V}$.
To estimate dense 3D DVF for landslide monitoring, we design a fusion-driven approach that jointly leverages TLS point clouds and co-registered RGB images. The RGB images, captured by built-in camera(s), are co-registered with the point cloud at the respective epoch (see details in Sec. 4.1). The proposed method consists of five key steps: tiling large point clouds and the corresponding RGB images (cf. Sec. 3.1); partitioning each point cloud tile into smaller patches (cf. Sec. 3.2); generating candidate coarse matches between patches (cf. Sec. 3.3); refining these coarse matches (cf. Sec. 3.4), and performing fine matching (cf. Sec. 3.5). An overview of the proposed method is illustrated in Fig. 1. While designed for two epochs, the method is easily generalizable to multi-epoch scenarios by sequentially processing pairs of epochs.
3.1 Point cloud and image tiling
Due to the large size of point clouds (e.g., tens of millions of points per epoch) representing an entire landslide, we divide the point clouds into smaller spatial tiles to improve computational efficiency and reduce memory usage. Following Gojcic et al. (2021), we first project the source and target point clouds to 2D along the coordinate axis that yields the largest bounding box in the projection. This bounding box is then recursively subdivided along the longer edge until each tile contains fewer than a predefined number of points (e.g., 1 million). Each tile pair, consisting of a source and a target tile that spatially overlap the same 3D region, is processed independently of the other tile pairs in the subsequent steps of our method. To accommodate displacements that may span tile boundaries, adjacent tiles are defined with an overlap margin equal to the maximum expected displacement (e.g., 10 m, set by a manually adjustable threshold). To ensure computational efficiency, the RGB images are divided into smaller overlapping tiles before being processed by the image matching algorithm.
3.2 Hierarchical point cloud partition
The local rigidity assumption is commonly applied in many applications (Sorkine and Alexa, 2007; Yang et al., 2014; Dewan et al., 2016). We argue that this assumption also holds for landslide cases, as small objects are more likely to move as rigid bodies. For example, individual boulders are expected to move as a single unit from one location to another, rather than splitting into separate pieces. To leverage this property, we propose partitioning the tiled point clouds into small patches, where patches correspond to those small objects or small regions undergoing locally consistent motion. Selecting an appropriate patch size is critical for accurate displacement estimation. If the patch size is too large, it may encompass multiple objects undergoing non-rigid motion, which breaks the local rigidity assumption. Conversely, patches that are too small are prone to containing mainly noise.
Figure 2: Hierarchical partitioning of the point cloud. Starting from the input point cloud colorized by RGB values, the method generates three levels of partitions that capture structural boundaries at multiple scales. Level i consists of the smallest and most numerous patches, while level iii contains the largest and fewest patches.
To capture multi-scale structural variations in the scene, we adopt the hierarchical partitioning strategy proposed by Robert et al. (2023), which decomposes the point cloud into patches at three levels of granularity (cf. Fig. 2) using pre-computed handcrafted features. We incorporate both geometric (linearity, planarity, and curvature) and radiometric (grayscaled RGB) features to guide this partitioning. Each partition level is used independently to guide patch-wise matching in the subsequent coarse-to-fine pipeline. At a later stage (cf. Sec. 3.5), we integrate the displacement estimates from all three levels to maintain high accuracy and high spatial coverage. Additionally, to mitigate the impact of noise, patches containing fewer than ten points are discarded for matching.
3.3 Candidate generation for coarse matching
For coarse matching, also known as patch matching, we define a match between a source and target patch as the pair with the highest similarity based on their geometric and radiometric attributes. A match may still be declared even if the patches (i) differ in their shape or (ii) undergo local orientation changes relative to their neighbors, as illustrated in Fig. 3. However, this matching procedure is impacted by the fact that the partitioning method rarely generates identical patches for the same regions across two epochs, even when the regions remain stable. To address this challenge and ensure accurate matching between patches from different epochs, we develop a processing scheme that is robust with respect to (w.r.t.) the exact patch shapes. This scheme contains patch matching derived from 3D source data (cf. Sec. 3.3.1) and patch matching derived from 2D source data (cf. Sec. 3.3.2). The final set of initial coarse matches combines both results, ensuring a balanced use of different types of information.
Figure 3: Different patch matching cases. Each case contains four example patch matches (same color) of point cloud $\mathbf{P}$ and $\mathbf{Q}$. Case (i): The patch match can differ in shape (see the dark cyan ones). Case (ii): The patch match can move differently compared to the neighbor patches (see the different orientations of the red arrows).
3.3.1 Patch matching from 3D source
Patch matching from the 3D source (i.e., point cloud) relies on 3D point feature matching. To ensure feature discriminativeness, we first downsample the point cloud tiles to achieve a uniform density. This prevents feature descriptors from over-focusing on densely populated regions while neglecting sparse areas. Unlike methods such as F2S3 (Gojcic et al., 2021)), which use a fixed voxel size for downsampling and feature extraction across all tiles, we adjust the voxel size adaptively based on the local density (i.e., the average scan resolution) of each tile. Next, we use a pre-trained local patch-based 3D neural network (Poiesi and Boscaini, 2021) to extract 3D point features from each point in the patch. Since the objective is to establish patch-wise matches, we then focus on aggregating the features of all points within each patch.
Figure 4: Coarse matching based on patch feature similarity. (i) Network architecture for constructing patch features using point features extracted from a pre-trained model within each patch. (ii) Mutual NN matching between patch features from the two epochs for generating candidate patch matches.
To achieve this, we design a lightweight network (cf. Fig. 4) and train it on publicly available indoor datasets (Zeng et al., 2017), eliminating the need for effort in annotating landslide training data. After the pre-trained model assigns features to the points within a patch, our network begins by using self-attention layers (Vaswani, 2017) to adaptively aggregate these features. The aggregated features then go through MLP layers to form the final patch features. Patches from the source and target point clouds are processed identically to construct patch features. During training, we apply the circle loss (Sun et al., 2020), a variant of the triplet loss (Choy et al., 2019) commonly used in 3D feature matching (Huang et al., 2021; Qin et al., 2022), to optimize the patch feature aggregation. In the testing phase, i.e., when estimating displacements for the real-world landslides, patch features are extracted using our pre-trained network. Patch-to-patch matches from the 3D source are then established through mutual NN matching based on their features.
3.3.2 Patch matching from 2D source
To ensure high spatial coverage of the output, we employ 2D pixel matching as a complementary approach. This approach operates on the full-resolution point clouds rather than the downsampled ones. Using the available camera intrinsic and extrinsic parameters, we first perform 2D pixel matching on the selected source and target RGB images. Then, we use a NN search to lift the established 2D pixel matches to 3D point matches, similar to the method we previously proposed in Wang et al. (2025a). However, point cloud tiles may span multiple RGB images, resulting in only a subset of points being validly projected onto any single RGB image. To address this issue, we associate each point cloud tile with its top-$k$ images, where the $k$-th image is the one with the $k$-th most projected points. Formally, for a given source point cloud tile $\mathbf{P}^{t}$, it is projected onto all available source images, and the top-$k$ images are selected. The same process is applied to the corresponding target point cloud tile. We apply image matching to the source and target images using a state-of-the-art deep-learning image matching algorithm, Efficient LoFTR (Wang et al., 2024). To integrate all 2D pixel matches across image pairs, we first select matches from the image pair with the most matches, adding matches from additional pairs only if the corresponding 3D points have not already been matched. Finally, we filter the lifted 3D point matches using a maximum displacement magnitude threshold, e.g., 10 m, to remove outliers with excessive displacement magnitudes. To construct patch matches, for each source patch, we count the number of matched target points falling within each target patch. The target patch with the highest count is considered the most match for the current source patch (see Fig. 5).
Figure 5: Coarse matching based on statistic analysis.
3.4 Patch matching refinement
The complete set of matches $\mathbf{M}$ extracted from both the 3D source and 2D source is given as:
$\mathbf{M}=\mathbf{M}_{3D}\cup\mathbf{M}_{2D}=\bigcup_{i=1}^{n}\{(\mathbf{p}_{i},\mathbf{q}_{i})\},$ (2)
where $\mathbf{M}_{3D}$ and $\mathbf{M}_{2D}$ represent patch matches established from the 3D and 2D source, respectively, and $p_{i}$ and $q_{i}$ $(i=1,...,n)$ are the 3D coordinates of the $n$ points matched in this way.
Figure 6: Patch matching refinement. To check the quality of a patch match. A MSDD criteria metric is computed that comparing the average pair distance between all pairs of two epochs, without the estimation of transformation.
We observe that mismatched patch matches will result in erroneous displacement estimates, as evidenced by the presence of outliers in the final displacement vectors. To mitigate this, we propose a refinement module that evaluates the quality of patch matches and filters out those with unreliable correspondences. Inspired by the as-isometric-as-possible constraint (Kilian et al., 2007; Huang et al., 2008) in computer graphics, our refinement module enforces local geometric consistency by preserving distances between points in the source and target patches. We introduce the Mean Squared Distance Deviation (MSDD) (cf. Eq. 3) as a measure of isometry preservation. A patch match is considered high quality if its MSDD is below a threshold $\delta_{1}$, and the proportion of squared distance deviations below $\delta_{1}$ exceeds a threshold $\delta_{2}$. These two thresholds are manually fine-tuned based on the specific characteristics of the dataset being processed. Only matches satisfying these criteria are retained for subsequent fine matching.
$\mathrm{MSDD}=\frac{1}{2N}\sum_{i=0}^{N-1}\sum_{j=0}^{N-1}\left|\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}-\|\mathbf{q}_{i}-\mathbf{q}_{j}\|^{2}\right|$ (3)
Here, $N$ denotes the number of correspondences (with $i\neq j$). A high MSDD value indicates significant deviation in the pairwise distances between corresponding points, reflecting distortion of their relative spatial structure and suggesting a poor-quality patch match. Conversely, a low MSDD value implies near-isometric consistency of structure, typically associated with a rigid transformation. By enforcing this criterion, we retain only geometrically consistent matches, thereby enhancing the reliability of displacement estimation.
3.5 Fine matching
Previous methods (Teza et al., 2007; Friedli and Wieser, 2016) estimate transformations by directly applying ICP between two-epoch point clouds or their tiles. However, we observe that this process often yields inaccurate results. This is because ICP and its variants (Rusinkiewicz and Levoy, 2001) rely on Euclidean distance comparisons, making them highly sensitive to noise and large transformations. To mitigate this issue, we incorporate an initial transformation estimation step that leverages our feature-based initial matches. Additionally, to ensure accurate estimation, even when matched patches exhibit diverse shapes, we utilize only point matches rather than all points within the matched patches.
Formally, for the patch matches $\mathbf{M}_{j}$, we employ the Kabsch algorithm (Kabsch, 1976) to estimate an initial rigid transformation $(\mathbf{R}_{j},\mathbf{T}_{j})$. This estimation is based on point matches $\{(\mathbf{p}_{i},\mathbf{q}_{i})\}_{i=1}^{n}$ within $\mathbf{M}_{j}$. This transformation is subsequently refined using point-to-point ICP (Besl and McKay, 1992), resulting in a refined transformation $(\hat{\mathbf{R}}_{j},\hat{\mathbf{T}}_{j})$. Although point matches are already relatively dense compared to most existing methods, to generate a more dense output, we apply the refined transformation to all source points $\mathbf{p}_{i}^{0}$ within the current patch match. We compute the 3D DVF restricted to the current patch j as:
$\mathbf{V}_{j}:\mathbf{p}_{i}^{0}\mapsto(\hat{\mathbf{R}}_{j}\cdot\mathbf{p}_{i}^{0}+\hat{\mathbf{T}}_{j}-\mathbf{p}_{i}^{0}),$ (4)
where $i$ ranges through all matched points in patch $j$ and $\mathbf{V}_{j}:(x,y,z)\mapsto(\Delta X,\Delta Y,\Delta Z)$ represents the per-point displacements in the current patch match. Finally, the 3D displacement vectors from all patch matches are collected to form the final 3D DVF:
$\mathbf{V}=[\mathbf{V}_{1},...,\mathbf{V}_{j},...,\mathbf{V}_{m}],$ (5)
Optionally, if exact point-wise correspondences are required, we further extract them based on the refined transformation estimated for each matched patch pair. For each transformed source point within a patch, we search for its nearest neighbor in the corresponding target patch. To reduce mismatches due to noise or structural variations, we discard correspondences where the point-wise distance exceeds a threshold (e.g., the average scan resolution).
As described in Section 3.2, we independently estimate displacements for each of the three hierarchical partition levels, followed by an integration step that consolidates the outputs. Level i, which offers the finest partitioning with the smallest and most numerous patches, is prioritized due to its strong consistency with the local rigidity assumption. However, to address cases of insufficient spatial coverage, we complement it with levels ii and iii. Specifically, for each point, we assign the displacement from level i if a valid estimate exists; otherwise, we fall back to level ii, and then to level iii.
4 Experiments
To evaluate the proposed method, we perform experiments on two large-scale landslide datasets. Section 4.1 provides details about study areas and datasets, while Section 4.2 and Section 4.3 outline the data preprocessing and evaluation steps, respectively.
4.1 Study areas and datasets
The first study area is located north of the Brienz village in Switzerland (see Fig. 7(a)). This area exhibits active slope movements with displacement rates reaching several meters per year (Krähenbühl and Nänni, 2017). These movements destabilize the upper rock mass, leading to large displacements and sporadic rockfalls. The Brienz TLS dataset (Kenner et al., 2022) was acquired using a Riegl VZ-6000 TLS scanner positioned near the road, approximately 1,500 m from the slope. The dataset captures landslide dynamics over an area of approximately 700 x 1000 m2. The scans have an average scan resolution of 0.08 m. RGB images were acquired using the built-in, calibrated 5-megapixel camera of the scanner. The camera captures images deflected by the laser mirror, fully covering the TLS field of view with dozens of overlapping images. Each image has a resolution of 2560 x 1920 pixels, a field of view of 7.2 x 5.5 (vertical x horizontal), a ground sampling distance (GSD) of approximately 0.05 m, and an exposure time of 5.5 ms. While manufacturer-supplied software (e.g., RiSCAN PRO v2.19.3 in our case) allows for automatic stitching into a panorama, we retain the individual images to preserve spatial accuracy and avoid stitching artifacts. For this study, we use TLS scans and their corresponding RGB images from two epochs: February and November of 2020. All images have the same intrinsic calibration but distinct poses relative to the scanner coordinate system, with parameters exported after the undistortion step in RiSCAN PRO. Additionally, observations from multiple total station (TS) prisms within the landslide area are used for quantitative evaluation.
Figure 7: Overview of study areas. (a): RGB image (2020) of the Brienz landslide area with the TLS scanner location indicated. (b): Map of Switzerland showing the locations of the two study areas; (c) RGB image (2019) of the Mattertal landslide area with the TLS scanner location. The two scanned landslide regions are highlighted using  green rounded rectangles. Base map source from ©Swisstopo.
The second study area is located in the Matter Valley in Switzerland (see Fig. 7(c)). This area contains various geologically active formations, including rock glaciers and landslides, with forested terrain in the lower part and rocky terrain at higher elevations (Möller et al., 2023; Shi et al., 2025). The area is subject to ongoing slope movements that could endanger local infrastructure and settlements. The Mattertal TLS dataset was acquired using a Riegl VZ-4000 TLS scanner positioned on stable bedrock on the opposite side of the valley, with a range of approximately 2,500 m. The dataset covers an area of approximately 3,000 x 2,000 m2. The scans have an average scan resolution of 0.30 m. For this study, we use TLS scans and RGB images from two epochs: July 2019 and July 2022. Similar to the Brienz dataset, we use RGB images captured by the built-in camera. These images have the same resolution as the Brienz dataset, but they have a GSD of approximately 0.20 m and an exposure time of 11 ms. We use RiSCAN PRO to undistort the images and export the intrinsic and pose information. Additionally, observations from multiple GNSS stations within the landslide area are used for quantitative evaluation.
4.2 Data preprocessing
The preprocessing steps include point cloud filtering and georeferencing. Point cloud filtering removes non-ground points, e.g., those from regions covered by vegetation. These points do not directly reflect the actual displacements of the landslide surface. We use CloudCompare v2.13.2 for this filtering process. The next step is georeferencing, which enables comparisons with external observations, e.g., measurements from TS prisms or GNSS stations. We implement georeferencing in two different ways according to the situation of each dataset. For the Brienz dataset, the landslide area is highly active, and a georeferencing matrix is available from the dataset provider for each epoch. Therefore, we simply apply the given transformation matrices to the two epoch scans for georeferencing. For the Mattertal dataset, however, most areas remain stable between analyzed epochs. We align the target epoch scan to the source epoch scan using RiSCAN PRO. We also derive a georeferencing matrix by aligning the source epoch scan to a Swiss digital elevation model using a global registration pipeline implemented in Open3D (Zhou et al., 2018). Georeferencing is achieved by applying the georeferencing matrix to the source epoch scan and aligned target epoch scans.
4.3 Evaluation
The evaluation of the proposed method consists of three aspects: (i) Accuracy of estimated displacement vectors is quantitatively assessed by comparison with external discrete observations, either from TS prisms or GNSS stations, and with manually picked reference points spatially closest to these observations. For each external observation, we identify the corresponding displacement vector by searching for its closest point in the source point cloud. This choice reflects the point-based nature of both the external observations and our outputs, enabling a direct one-to-one correspondence that avoids potential spatial smoothing. The accuracy of manual picking aligns with the average scan resolution. While alternative strategies ((Gojcic et al., 2021; Shi et al., 2025; Wang et al., 2025a) involving spatial averaging may improve robustness to noise, they risk diluting localized displacement signals and are therefore deferred to A for supplementary analysis. (ii) Spatial coverage is quantified by calculating the ratio of valid displacement estimates to the total number of points in the source point cloud for selected regions of interest (ROIs). To prevent counting duplicated estimates, both displacement estimates and source point clouds are downsampled using a voxel size corresponding to the average scan resolution. (iii) Qualitative evaluation is performed to assess the spatial distribution and directionality of the estimated displacement vectors for consistency with established geomorphological understanding.
5 Results and discussion
In this section, we first introduce the baseline methods used for comparison in 3D displacement estimation in Section 5.1. We then present the accuracy evaluation in Section 5.2, followed by the spatial coverage evaluation in Section 5.3. The effectiveness of the proposed fusion strategy is examined in Section 5.4, along with the discussion of the impact of multi-level partitioning in Section 5.5. We further assess the contribution of the refinement module in Section 5.6. Finally, the limitations of the proposed method and potential improvements are discussed in Section 5.7.
5.1 Baseline selection
We compare our method against three existing approaches for estimating dense 3D DVFs in landslide monitoring: Piecewise ICP (Friedli and Wieser, 2016), F2S3 (Gojcic et al., 2021), and our previously proposed RGB-Guided method (Wang et al., 2025a). These methods serve as representatives of Categories II, III and IV, as introduced in Section 2.1. Categories I and V are omitted, as the former does not yield full 3D displacement vectors, and the latter is not designed for TLS point clouds. To enable full-scene landslide analysis, we adapt the Piecewise ICP and RGB-Guided methods, which are limited to single regions of interest, by incorporating the tiling process. For other settings, we follow the choices described in the respective publications.
5.2 Accuracy evaluation
Tab. 1 shows the quantitative results of our method for both landslides. We compare the estimated 3D displacements of our method with external observations: TS prisms on the Brienz dataset and GNSS stations on the Mattertal dataset. For clarity of interpretation, the coordinate system for both datasets is defined following the right-hand rule, where the Y-axis approximately corresponds to the horizontal projection of the mean line-of-sight (LOS) direction, the X-axis is perpendicular to the Y-axis horizontally, and the Z-axis points vertically upward. The displacement magnitude deviations (i.e., $|\Delta D_{S}|$) between our estimates and the external observations are 0.15 m and 0.25 m, respectively. They are lower than the corresponding average footprint sizes of 0.23 m and 0.68 m, respectively. Additionally, we manually select corresponding points at the second epoch (i.e., target epoch) to derive displacements that accurately reflect the input data used in both our method and other baseline approaches. The deviations between our method and the manually selected reference displacements are 0.07 m and 0.20 m for the Brienz and Mattertal datasets, respectively. These deviations are smaller than the average scan resolutions of 0.08 m and 0.30 m for these two datasets, respectively. This demonstrates the high accuracy of the displacement estimation by our method. Furthermore, our method has larger deviations with external observations than with manual reference data, which can be attributed to georeferencing errors. To avoid this, we use the manually picked points as the reference for comparing our method with baseline methods.
Table 1: Comparison of 3D displacement estimates from our method (based on NN search) against external observations and manually labeled data. The upper part of the table shows results for the Brienz landslide, evaluated using TS observations and manually labeled data; the lower part presents results for the Mattertal landslide, evaluated using GNSS observations and manually labeled data. $|\Delta D_{X}|$, $|\Delta D_{Y}|$, $|\Delta D_{Z}|$ denote the absolute deviations in the individual displacement components, while $|\Delta D_{S}|$ represents the absolute deviation in the displacement magnitude.
Ours (m) $|\mathrm{Ours-External}|$ (m) $|\mathrm{Ours-Manual}|$ (m)
Point Range (km) Footprint (m) $D_{X}$ $D_{Y}$ $D_{Z}$ $D_{S}$ $|\Delta D_{X}|$ $|\Delta D_{Y}|$ $|\Delta D_{Z}|$ $|\Delta D_{S}|$ $|\Delta D_{X}|$ $|\Delta D_{Y}|$ $|\Delta D_{Z}|$ $|\Delta D_{S}|$
Quantitative results on the Brienz dataset
T1 1.97 0.24  0.36 -1.84 -4.36 4.75 0.16 0.04 0.06 0.19 0.09 0.08 0.11 0.13
T2 2.00 0.24  0.11 -0.21 -0.30 0.38 0.07 0.03 0.17 0.23 0.08 0.15 0.10 0.17
T3 1.65 0.20  0.24 -0.38 -0.64 0.78 0.46 0.05 0.01 0.00 0.06 0.15 0.28 0.12
T4 1.80 0.22  0.31 -0.68 -0.60 0.96 0.13 0.01 0.36 0.28 0.01 0.08 0.06 0.02
T5 1.91 0.23  0.91 -0.94 -1.23 1.80 0.07 0.05 0.26 0.05 0.17 0.03 0.03 0.05
T6 1.81 0.22  0.55 -0.82 -0.70 1.21 0.07 0.00 0.29 0.16 0.09 0.14 0.20 0.03
T7 1.88 0.23  1.13 -1.54 -1.94 2.72 0.01 0.07 0.09 0.16 0.01 0.04 0.04 0.00
Average 1.90 0.23  -  -  - - 0.14 0.04 0.17 0.15 0.07 0.09 0.12 0.07
Quantitative results on the Mattertal dataset
G1 4.44 0.67 -0.06 -0.08 -0.17 0.20 0.06 0.08 0.18 0.19 0.01 0.02 0.12 0.07
G2 4.58 0.69 -3.00 -4.39 -1.90 5.65 0.94 0.01 0.27 0.52 0.48 0.06 0.02 0.28
G3 4.62 0.69  0.95 -1.01 -0.50 1.47 0.50 0.06 0.12 0.35 0.48 0.06 0.19 0.26
G4 4.40 0.66  0.29 -0.83 -0.30 0.93 0.03 0.14 0.01 0.13 0.16 0.07 0.41 0.21
Average 4.51 0.68  -  -  - - 0.37 0.05 0.10 0.25 0.28 0.05 0.19 0.20
Figure 8: Quantitative comparison of estimated displacements from different methods w.r.t. the manual reference, focusing on deviations along the X- and Y-axis directions (i.e., $\Delta D_{X}$ and $\Delta D_{Y}$). Dashed circles and grid lines are added to improve readability. Similar deviation patterns are observed along the X- and Z-axis directions as well (see B).
We further compare the estimated displacements of our method with those of baseline methods in Fig. 8, where we use manually picked points as the reference and compare the deviations between the different methods and the reference. Our method shows maximum deviations of 0.17 m along the X-axis for the Brienz dataset and 0.48 m along the Y-axis for the Mattertal dataset. Among all baseline methods, F2S3 achieves the closest performance to ours, with maximum deviations of 0.18 m along the Y-axis for the Brienz dataset and 0.57 m along the Y-axis for the Mattertal dataset. It is worth noting that all baseline methods fail to provide valid estimates corresponding to all external observations (e.g., Piecewise ICP only provides two valid estimates for all four GNSS stations on the Mattertal dataset) due to their reliance on only one type of information (either 3D point cloud or 2D RGB information). In contrast, our method provides estimates around all external observations. Piecewise ICP (Friedli and Wieser, 2016) exhibits the largest deviations on both datasets, primarily due to its approach of uniformly partitioning and subsequently matching the closest patches based on Euclidean distance. The RGB-Guided approach shows large deviations along the Y-axis direction. An investigation in Wang et al. (2023) suggests that, even though the manufacturer provides co-registration between the TLS scanner and its built-in camera(s), misalignment errors of several pixels can still occur between the point clouds and the images. These errors persist across epochs due to variations in the camera poses associated with the captured images, even when the same scanner is used. This is particularly problematic on the Mattertal dataset, where the average GSD is 0.2 m, resulting in potential meter-level displacement errors. In our approach, 2D information mainly contributes to patch matching. Due to the refinement, our method with fused 3D and 2D information still is comparable in accuracy to the 3D-based approach, F2S3.
5.3 Spatial coverage evaluation
We present the 3D displacements estimated by our method in Fig. 9 and Fig. 10 for the Brienz and Mattertal datasets, respectively. These results demonstrate that our method can estimate 3D displacements across most landslide areas, with spatial coverage of 79% and 97% w.r.t. the source point cloud for the Brienz and Mattertal datasets, respectively. The Brienz landslide exhibits significant variation in displacement magnitude across the area, whereas the Mattertal landslide remains relatively stable, with only a few localized areas showing notable movement between the observed epochs. To further analyze the displacement characteristics, we select a ROI from the highly active areas of each landslide, where several external observations are also included in the corresponding ROI. In the Brienz ROI (see Fig. 9), the 3D displacement vectors estimated by our method have a mean magnitude of approximately 2.5 m and mean azimuth and elevation angles of approximately 135° and -45°, respectively. This suggests displacement toward the southeast. In contrast, most of the displacement magnitudes in the Mattertal landslide ROI (see Fig. 10) are close to zero, and the azimuth and elevation angles exhibit large spreads without consistent direction trends. This indicates a lack of coherent surface motion during the observed period.
Figure 9: Estimated 3D displacements on the Brienz landslide. (a): Displacement magnitudes across the entire landslide (ROI highlighted in  red). (b) Azimuth direction of displacements within the ROI. (c): Elevation direction of displacements within the ROI. The source point cloud is shown in gray as the background for better readability, which applies to all other figures in this section.
Figure 10: Estimated 3D displacements on the Mattertal landslide. (a): Displacement magnitudes across the entire landslide (ROI highlighted in  red). (b) Azimuth direction of displacements within the ROI. (c): Elevation direction of displacements within the ROI.
Fig. 11 and Fig. 12 present qualitative comparisons between our method and baseline methods for the ROIs of the Brienz and Mattertal datasets, respectively. While Piecewise ICP achieves high spatial coverage (93% and 95%) on both ROIs, it underestimates the actual displacements on the Brienz dataset and produces outlier estimates on the Mattertal dataset. In contrast, the other three methods produce results that better reflect the actual displacement trends. For both ROIs, F2S3 achieves spatial coverage of 38% and 52%, and the RGB-Guided approach achieves spatial coverage of 45% and 66%. These are both much lower than the coverage of 82% and 90% achieved by our method. This is attributed to the partition-based coarse-to-fine estimation and the integrated use of both 3D and 2D information in our method. Furthermore, the RGB-Guided approach produces many noisy estimates on the Mattertal dataset, primarily due to the aforementioned co-registration misalignment errors (see Sec. 5.2). In our method, noisy correspondences from the 2D source are largely suppressed due to the adaptive balance between 3D and 2D information.
Figure 11: Qualitative comparison of estimated displacement magnitudes using different methods on the ROI of the Brienz landslide.
Figure 12: Qualitative comparison of estimated displacement magnitudes using different methods on the ROI of the Mattertal landslide.
5.4 Effectiveness of the fusion module
We evaluate the effectiveness of the fusion module, with qualitative results shown in Fig. 13. For comparison, we define two ablated variants: Ours-2D, which performs both coarse and fine matching using only RGB image information, and Ours-3D, which relies solely on geometric information from the point cloud. We observe that the RGB-Guided method overestimates the displacement magnitudes, consistent with quantitative results in Sec. 5.2. This may be attributed to the co-registration error, which also decreases the quality of 3D displacements estimated by Ours-2D. Conversely, Ours-3D yields more accurate estimates but fails in regions dominated by debris, where geometric features are less discriminative. By integrating both 2D RGB and 3D geometry through the fusion module, our method improves spatial coverage to 90% compared to 87% of Ours-3D, while mitigating noisy correspondences introduced by 2D RGB matching alone.
Figure 13: Comparison of our method using only 2D information, only 3D information, and both 2D and 3D information on the ROI of the Mattertal landslide.
5.5 Impact of multiple levels of partitions
We analyze the impact of using different partition strategies, with qualitative results shown in Fig. 14. We apply each of our three hierarchical levels (level i, ii, iii) of partitions independently using the same matching pipeline. The supervoxel-based partition adopted in F2S3 serves as a baseline and represents a single-level approach. It produces displacements with 74% spatial coverage, which is higher than our level i (55%) but lower than level ii (79%). While level iii offers higher spatial coverage (90%), it tends to over-smooth displacements by breaking object boundaries. In contrast, our proposed hierarchical method (cf. ours) integrates all three levels incrementally, yielding 90% coverage while better preserving local structural details.
Figure 14: Comparison of our method using different partitioning strategies or levels on the ROI of the Mattertal landslide.
5.6 Effectiveness of the refinement module
We analyze the impact of the refinement module and present the qualitative results in Fig. 15. We compare three cases: (i) The original 3D result obtained directly from the pre-trained deep learning model; (ii) the result obtained using our method without the refinement module; and (iii) the result obtained using our method with the refinement module. The original 3D result has relatively low spatial coverage (54%) and many noisy estimates, which can be attributed to two factors: 1) The 3D matching fails to produce a sufficient number of matches due to the limitations of geometric features in some regions, where RGB features could contribute, and 2) the absence of the refinement module. When we compare results with and without the refinement module, we find that our method without refinement yields an output with higher spatial coverage (92%) but more noisy estimates. In contrast, our method with the refinement module effectively eliminates low-quality patch matches, primarily consisting of noisy correspondences, by applying a strict and practical filtering criterion. This improves overall match reliability while maintaining high spatial coverage (90%) of the output.
Figure 15: Comparison of estimated displacement magnitudes: original 3D output, our method without the refinement module (cf. Ours w/o refine.), and our method with the refinement module (cf. Ours) on the ROI of the Mattertal landslide.
5.7 Limitations and potential improvements
Limitations: Our approach has one distinct limitation: it relies on feature distinctiveness for accurate point-wise or pixel-wise matching. In regions with repetitive geometric structures or weak radiometric texture, e.g., debris-covered surfaces, or in cases of severe occlusion and topographic disruption caused by large toppling or rotational failures, the method may fail to establish accurate correspondences, resulting in missing displacement estimates. Potential improvements: (i) We adopt the 3D patch-based model from (Poiesi and Boscaini, 2021) and the image matching algorithm from (Wang et al., 2024), both of which are commonly used in related applications. While they currently provide strong performance, we believe that future advances in feature learning and image matching may further improve the spatial coverage and accuracy. (ii) The method is applied to TLS datasets with images captured by built-in, calibrated cameras. Since these cameras are closely aligned with the scanner in terms of viewpoint, fusion benefits are limited to the feature level. We anticipate that integrating images from different viewpoints, such as stationary cameras or UAV platforms, could introduce complementary information at the data level. We leave this exploration for future work.
6 Conclusion
This paper presents a hierarchical partition-based approach for estimating dense 3D DVF in TLS-based landslide monitoring. The method optimizes the matching pipeline in a coarse-to-fine manner using both 3D point clouds and 2D RGB images. By integrating multi-level partition information, we effectively capture structural variations and mitigate the occurrence of noisy estimates while maintaining high spatial coverage. Furthermore, rather than requiring annotated landslide-specific training datasets, which are hard to obtain, the method leverages features extracted from models pre-trained on publicly available indoor datasets. Experiments on two real-world landslide datasets with different displacement patterns validate the effectiveness of the proposed method: it produces displacement estimates with higher spatial coverage than the current state-of-the-art method (F2S3) while maintaining comparable accuracy. Although this method has been verified for TLS-based landslide monitoring, it can be extended to other types of point clouds and other monitoring applications. Future research will focus on exploring the integration of TLS point clouds with RGB images captured from external cameras, such as those mounted on UAVs. This could provide multi-perspective data, enabling landslide monitoring with an enhanced spatial level and coverage.
CRediT authorship contribution statement
Zhaoyi Wang: Conceptualization, Methodology, Data curation, Investigation, Formal analysis, Validation, Visualization, Writing – original draft, Writing – review & editing, Software. Jemil Avers Butt: Conceptualization, Methodology, Formal analysis, Writing – review & editing, Supervision. Shengyu Huang: Conceptualization, Methodology. Tomislav Medić: Methodology, Writing - Review & Editing. Andreas Wieser: Conceptualization, Writing – review & editing, Resources, Supervision, Project administration, Funding acquisition.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to assist in refining the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
The Brienz dataset was provided by Robert Kenner (Institute for Snow and Avalanche Research), and total station observations were provided by Stefan Schneider (CSD INGENIEURE AG). The data collection of the Mattertal dataset was financially supported by the Federal Office for the Environment (FOEN) and by the Swiss Geodetic Commission (SGC). Helena Laasch has shared the code for Piecewise ICP. The first author was supported by the China Scholarship Council (CSC).
Appendix A Comparison with averaged displacements
Following (Gojcic et al., 2021; Shi et al., 2025; Wang et al., 2025a), for each external observation or manually pickled point, we select a set of estimated displacement vectors within a certain radius. We use a 15 m radius for both datasets in default. However, for point T5, we observe a significant deviation in the mean absolute deviation (MAD) due to local non-rigid motion. Therefore, we reduce the radius to 2.5 m for point T5. The average deviations between the estimated displacement magnitudes of our method and the external observations are 0.12 m and 0.30 m for the two datasets, respectively, both smaller than the corresponding footprint sizes of 0.23 m and 0.68 m. The average deviations between estimated displacement magnitudes of our method and external observations are 0.10 m and 0.12 m, both smaller than the average scan resolutions of 0.08 m and 0.30 m. Theses findings are consistent with our analysis based NN search in Section 5.2. The MAD of the estimated displacement magnitudes are 0.15 m and 0.24 m, respectively, suggesting the high precision of the displacement magnitudes estimated by our method.
Table 2: Comparison of 3D displacement estimates from our method (based on averaging) against external observations and manually labeled data. The upper part of the table shows results for the Brienz landslide, evaluated using TS observations and manually labeled data; the lower part presents results for the Mattertal landslide, evaluated using GNSS observations and manually labeled data. $\overline{D_{X}}$, $\overline{D_{Y}}$, $\overline{D_{Z}}$, denote the mean estimated displacement components, and $\overline{D_{S}}$ denotes the mean estimated displacement magnitude
Ours (m) $|\mathrm{Ours-External}|$ (m) $|\mathrm{Ours-Manual}|$ (m) MAD (m)
Point Range (km) Footprint (m) $\overline{D_{X}}$ $\overline{D_{Y}}$ $\overline{D_{Z}}$ $\overline{D_{S}}$ $|\Delta D_{X}|$ $|\Delta D_{Y}|$ $|\Delta D_{Z}|$ $|\Delta D_{S}|$ $|\Delta D_{X}|$ $|\Delta D_{Y}|$ $|\Delta D_{Z}|$ $|\Delta D_{S}|$ $\mathrm{MAD}(D_{X})$ $\mathrm{MAD}(D_{Y})$ $\mathrm{MAD}(D_{Z})$ $\mathrm{MAD}(D_{S})$
Quantitative results on the Brienz dataset
T1 1.97 0.24 -0.08 -1.92 -4.09 4.65 0.28 0.12 0.32 0.10 0.35 0.00 0.38 0.22 0.53 0.27 0.38 0.30
T2 2.00 0.24  0.18 -0.17 -0.30 0.41 0.15 0.00 0.17 0.20 0.16 0.11 0.09 0.20 0.04 0.04 0.05 0.04
T3 1.65 0.20 -0.23 -0.39 -0.56 0.78 0.93 0.06 0.08 0.01 0.42 0.14 0.20 0.12 0.58 0.13 0.15 0.22
T4 1.80 0.22  0.06 -0.69 -0.67 1.00 0.38 0.02 0.28 0.24 0.23 0.09 0.02 0.06 0.25 0.12 0.10 0.11
T5 1.91 0.23  0.94 -0.92 -1.25 1.80 0.05 0.03 0.23 0.05 0.14 0.01 0.05 0.05 0.06 0.03 0.03 0.04
T6 1.81 0.22  0.08 -0.86 -0.62 1.25 0.54 0.04 0.37 0.12 0.38 0.10 0.12 0.07 0.45 0.31 0.17 0.29
T7 1.88 0.23  1.15 -1.47 -1.92 2.71 0.02 0.00 0.11 0.15 0.02 0.10 0.01 0.01 0.06 0.10 0.09 0.08
Average 1.86 0.23  -  -  - - 0.34 0.04 0.22 0.12 0.24 0.08 0.12 0.10 0.28 0.14 0.14 0.15
Quantitative results on the Mattertal dataset
G1 4.44 0.67 -0.07 -0.09 -0.15 0.17 0.07 0.09 0.16 0.16 0.02 0.01 0.10 0.04 0.03 0.02 0.03 0.04
G2 4.58 0.69 -2.60 -4.22 -1.77 5.73 0.54 0.19 0.14 0.61 0.08 0.12 0.15 0.37 0.51 0.18 0.14 0.23
G3 4.62 0.69  0.94 -0.98 -0.48 1.24 0.49 0.02 0.11 0.12 0.47 0.10 0.18 0.03 0.09 0.13 0.09 0.27
G4 4.40 0.66  0.44 -0.72 -0.35 1.11 0.18 0.03 0.03 0.30 0.01 0.04 0.37 0.04 0.50 0.30 0.30 0.40
Average 4.51 0.68  -  -  - - 0.32 0.08 0.11 0.30 0.15 0.07 0.20 0.12 0.28 0.16 0.14 0.24
Appendix B Deviation patterns along the X- and Z-axis directions
Figure 16: Quantitative comparison of estimated displacements from different methods w.r.t. the manual reference, focusing on deviations along the X- and Z-axis directions (i.e., $\Delta D_{X}$ and $\Delta D_{Z}$). Dashed circles and grid lines are added to improve readability.